# Fast Polynomial Root-Finding and Approximation of the Eigenvalues of a Matrix

### v36

@@ 14.5 Counting the evaluation queries of our root-finders
Combine Cors. 2.2 and 14.1 and Prop. 14.4 and obtain
@@
We can reduce root-finding to deterministic evaluation of $p(x)$ at $O(m^{2}b\log(d))$ points as well as to evaluation at $O(mb\log^{2}(d))$ points under the Random Root Model.
15 Reduction of $m$-test to e/i tests
Our two algorithms of Secs. 15.1 and 15.2 reduce $m$-test for a disc $D$ to bounded number of e/i tests. Namely, the algorithm of Sec. 15.2 uses just a single e/i test where $m=d$ but up to order of $\log(d)$ e/i tests otherwise, while the algorithm of Sec. 15.1 uses $O(1)$ e/i tests for any $m$.
15.1 $m$-test by means of e/i tests for any disc
Our alternative reduction of $m$-test for $1\leq m\leq d$ to a bounded number of e/i tests can be applied in the case of any disc, not necessarily isolated. If at least one of the e/i tests fails, then definitely there is a root in a narrow annulus about the boundary circle of an input disc, and hence the internal disc of the annulus would not contain $m$ roots. Otherwise we would obtain a root-free concentric annulus about the boundary circle of an input disc and then reduce $m$-test to root-counting in an isolated disc, which we can perform at a low cost (see, e.g., claim (ii) of Prop. 7.1).
Algorithm 15.1.
Reduction of $m$-test to e/i tests.
INPUT: a polynomial $p(x)$ of a degree $d$, two values $\bar{\theta}>1$ and $\bar{\sigma}>1$, an integer $m$ such that $1\leq m\leq d$ and #$\#(D(0,\bar{\sigma}))\leq m$, and two algorithms $\mathbb{ALG}~{}0$ and $\mathbb{ALG}~{}1$, which for a black box polynomial perform $\bar{\sigma}$-soft e/i test for in any disc and root-counting in an isolated disc, respectively.
$v=\lceil 2\pi/\phi\rceil$ (15.1)
$c_{j}=\exp(\phi_{j}{\bf i}),~{}j=0,1,\dots,v,$ (15.2)
$c_{1}$$c_{2}$$c_{3}$$c_{4}$$c_{5}$$c_{6}$$c_{7}$$c_{8}$$c_{9}$$c_{0}$
Figure 6: $v=9$, $|c_{j}-c_{j-1}|=\rho>|c_{9}-c_{0}|$ for $j=1,2,\dots,9$.
To prove correctness of Alg. 15.1 we need the following result.
Theorem 15.1.
Proof.
$c_{0}$$c_{1}$$z$$z^{\prime}$
Let us prove that $|z_{j}-z_{j}^{\prime}|=\rho\sqrt{3}$ for $j=1$, and hence also for $j=2,\dots,v$.
Corollary 15.1.
Correctness of Alg. 15.1. (i) If algorithm $\mathbb{ALG}~{}1$ outputs inclusion in at least one of its applications, then one of the discs $D(c_{j},\bar{\sigma}\rho)$ contains a root, and then Alg. 15.1 correctly reports that $r_{d-m+1}\geq 1-\sigma\bar{\rho}$.
(ii) Otherwise Cor. 15.1 holds, and then correctness of Alg. 15.1 follows from correctness of $\mathbb{ALG}~{}1$.
Computational complexity. Alg. 15.1 invokes algorithm $\mathbb{ALG}~{}0$ at most $v+1$ times for $v=O(1)$ of (15.1). If $\mathbb{ALG}~{}0$ outputs exclusion $v+1$ times, then algorithm $\mathbb{ALG}~{}1$ is invoked and runs at a dominated cost.
Remark 15.1.
One can modify Alg. 15.1 as follows (see [63]): instead of fixing a positive $\rho$ and then computing $v$ of (15.1) fix an integer $v>0$, apply algorithm $\mathbb{ALG}~{}0$ at $v$ equally spaced points $c_{j}=\exp(\frac{2j\pi{\bf i}}{v}),~{}j=0,1,\dots,v-1$, let $\phi=\frac{2\pi}{v}$, and then define $\rho$. We also arrive at this algorithm if we compute or guess this value $\rho$ and then apply Alg. 15.1, but notice that $c_{v}=c_{0}$ for $c_{j}$ of (15.2) and only invoke algorithm $\mathbb{ALG}~{}0$ for $c_{0},\dots,c_{v-1}$.
15.2 Simple reduction to a single e/i test in the case of an isolated disc
Clearly, we can immediately reduce the $d$-test in the unit disc $D(0,1)$ for $p(x)$ to performing 1-test (that is, e/i test) in that disc for the reverse polynomial $p_{\rm rev}(x)$ of (1.5).
Let us next reduce an $m$-test, for $0<m<d$, to a single application of 1-test provided that an input disc $D$ with $m$ roots is $\theta$-isolated for $\theta$ of order $\log(d)$; we can ensure that provision by applying root-lifting or root-squaring, in both cases at the cost of evaluation of $p(x)$ at $O(\log(d))$ points if an initial disc is just isolated.
Proposition 15.1.
Let the root sets of two polynomials $f(y)$ and $g(y)$ of degrees $m$ and $d-m$, respectively, lie outside the disc $D(0,1)$ and in the disc $D(0,1/\theta)$, respectively, for a sufficiently large real $\theta$. Then
${\rm NIR}_{f}(z)={\rm NIR}_{fg}(z)-\frac{d-m}{z}+\Delta~{}{\rm for}~{}|z|=1~{}{\rm and}~{}|\Delta|\leq\frac{d-m}{\theta-1}.$
Proof.
Notice that $\frac{1}{z-y_{j}}-\frac{1}{z}=\frac{y_{j}}{(z-y_{j})z}$, recall that $|y_{j}|\leq\frac{1}{\theta}$ for $\theta>1$ and all $d-m$ roots $y_{1},\dots,y_{d-m}$ of $g(y)$ by assumption, and deduce that
$\Big{|}\frac{1}{z-y_{j}}-\frac{1}{z}\Big{|}\leq\frac{1}{\theta-1}~{}{\rm for}~{}|z|=1~{}{\rm and}~{}j=1,\dots,d-m.$ (15.4)
[…]
Next fix an integer $q$ such that $m<q\leq 2m$, e.g., $q:=m+1$ or $q:=2^{h}$ for $h=\lceil\log_{2}(m+1)$. Assume that $\theta-1$ exceeds a positive constant and ensure that
$\theta>1+4(d-m)\sqrt{q}$ (15.5)
by applying root-lifting or root-squaring at the cost of evaluation of $p(x)$ at $O(\log(d))$ points.
[…]
$V_{f}:=\max_{g=0}^{q-1}|{\rm NIR}_{fg}(\zeta^{g})-(d-m)\zeta^{-g}|.$ (15.6)
Now ensure that $V_{f}<\frac{1}{2\sqrt{q}}$, say, by requiring that $\max_{g=0}^{q-1}|{\rm NIR}_{fg}(\zeta^{g})|<\frac{1}{4\sqrt{q}}$ because bound (15.5) implies that $\frac{d-m}{\theta-1}<\frac{1}{4\sqrt{q}}$.
@@ 16.2 A priori estimates for computational precision
In the rest of this section we deduce a priori estimate of $O(b\log(d))$ bits for computational precision, proving that this precision supports correctness of our subdivision root-finders. Such an estimate is deduced in [76] based on advanced non-trivial techniques, of which we only use a very small part.
Estimating computational precision of our soft e/i tests we assume that they are applied to the unit disc $D(0,1)$ because we can extend the results to any disc $D(c,\rho)$ by applying map (1.3) and increasing the upper bounds on the precision by a factor of $\log_{2}(1/\rho)$. This factor is at most $b$ because we only handle discs of radii $\rho>\epsilon=1/2^{b}$ until we stop reaching $\epsilon$-discs. Therefore, our bound $O(\log(d))$ on the computational precision of our e/i tests for the disc $D(0,1)$ will imply the bound $O(b\log(d))$ for any disc.
We also make Assumption 0 that we can request and obtain from evaluation oracle (black box subroutine) the values of $p$ with a relative error bound in $1/d^{\nu}$ for any fixed constant $\nu>0$ of our choice; we can represent these values with the bit-precision $\log_{2}(d)+\log_{2}(\nu)$.
In our e/i tests we only need to compare the values $|{\rm NIR}_{t_{h}}(x)|$ at some complex points $x$ with some fixed values $v_{d}\geq 1/d^{\beta}$ for a constant $\beta>0$, and next we estimate that under Assumption 0 we can indeed do this by performing our computations with a precision of $O(\log(d))$ bits.
Multiplication with a precision $s$ contributes a relative error at most $2^{s}$ (cf. [21, Eqn. A.3 of Ch. 3]) and otherwise just sums the relative error bounds of the factors ignoring the dominated impact of higher order errors. Hence we deduce that $h=O(\log(d))$ multiplications for root-lifting to their $h$th powers increase these bounds by at most a factor of $h$, which is immaterial since we allow relative errors of order $1/d^{\nu}$ for a constant $\nu>0$ of our choice.
We only apply this analysis to approximation of NIR$(x)$ for $p_{h}(x)$ based on (3.2) and also use the following
@@
This assumption holds because we apply our e/i tests under random rotation $x\mapsto vx$ for the random variable $v$ specified in our next proposition.
@@
Let $\psi_{j}$ denote the length of the arc $C(0,1)\cap D(z_{j},\frac{2\pi}{mqw})$, $j=1,2,\dots,d$. Then $\psi_{j}=0$ for at least $d-m$ integers $j$, and $\max_{j=1}^{d}\psi_{j}\leq\frac{\pi^{2}}{mqw}$ for sure. Hence $\sum_{j=1}^{d}\psi_{j}<\frac{\pi^{2}}{qw}$, and so $\phi_{g}\leq\frac{2\pi}{mqw}$ for a fixed $g$ with a probability at most $\frac{\pi}{qw}$. Therefore, $\phi\leq\frac{2\pi}{mqw}$ with a probability at most $\frac{\pi}{w}$. ∎
Next we prove that a precision of $O(\log(d))$ bits supports our e/i tests under Assumptions 0 and 1 by applying Prop. 16.1 for sufficiently large $w=d^{O(1)}$. We will (i) prove that $|{\rm NIR}(z)|=d^{O(1)}$, (ii) approximate this value within $d^{\nu}\delta$ for any fixed constant $\nu$ and for $\delta$ of (3.2) chosen such that $\log(\frac{1}{\delta})=O(\log(d))$, and (iii) estimate that such a bound withstands the impact of rounding errors of computations for (3.2).
@@
In the proof of the next theorem we use the following lemma, which is [76, Fact 3.5].
@@
Under the assumptions of Prop. 16.2 it holds that
$|t(x)\alpha|=|\frac{t(x)-t(x-\delta)}{\delta}-t^{\prime}(x)|\leq|t(x)\frac{(d\xi^{\prime})^{2}\delta}{1-d\xi^{\prime}\delta}|$ for $\xi^{\prime}=\frac{1}{\rho}$.
@@
The claimed bound on $|\xi^{\prime}|$ follows from (3.1). It remains to apply the first five lines of the proof of [76, Lemma 3.6] with $f$, $\alpha$, $n$, and $\xi$ replaced by $t$, $-\delta$, $d$, and $w$, respectively. Namely, first obtain from Taylor’s expansion that $t(x)-t(x-\delta)=\sum_{j=0}^{\infty}\frac{\delta^{j}}{j!}t^{(j)}(x)-t(x)$.
@@
Suppose that the values $t(y)$ have been computed by a black box oracle within a relative error bounds $\nabla(y)$ for $y=x$, $y=x-\delta$, and $\delta$ of Cor. 16.2. Then one can ensure that $\beta<\frac{1}{d^{\bar{\eta}}}$ for any fixed constant $\bar{\eta}$ by choosing a proper $\nabla(y)$ of order $1/d^{O(1)}$, represented with a precision of $O(\log(d))$ bits, which we can do under Assumption 0.
@@
$\frac{t(x-\delta)}{t(x)}=1+\frac{t(x-\delta)-t(x)}{t(x)}=1+\delta\frac{t^{\prime}(u)}{t(x)}$, for $u\in[x-\delta,x]$.
@@
By extending Cor. 16.2 obtain $|\alpha^{\prime}|\leq\frac{1}{d^{O(1)}}$ and deduce from (3.1) and Assumptions 0 and 1 that $|\frac{t^{\prime}(x)}{t(x)}|\leq\frac{1}{d^{O(1)}}$ (cf. the proof of Prop. 10.1). Hence $w:=|\frac{t(x-\delta)}{t(x)}|\leq d^{O(1)}$ and $|\beta\delta|\leq|w|\frac{|\nabla(x)|+|\nabla(x-\delta)|}{1-|\nabla(x)|}$.
Now choose $\nabla(y)$ such that $|\nabla(y)|\leq\frac{|\delta|}{20|w|}$ for $y=x$ and $y=x-\delta$. Then verify that $\log(\frac{1}{|\nabla(y)|})=O(\log(d))$, $1-|\nabla(x)|\geq 1-\frac{|\delta|}{20|w|}>\frac{19}{20}$, and $|\delta\beta|<\frac{|\delta|}{8}$. ∎
Combine Cors. 16.1 and 16.2 and Prop. 16.4 to obtain
@@
Given a black box (oracle) for the evaluation of a polynomial $p(x)$ of a degree $d$ with a relative error of any order in $1/d^{O(1)}$ (this holds under Assumption 0), we can perform our black box subdivision root-finders with a precision of $O(\log(d))$ bits.
@@
We derived our precision bound $O(\log(d))$ (in the case of root-finding in the unit disc $D(0,1)$) based on operating entirely with NIR$(x)$. For comparison, if we relied on computing $p(x)$ and on the bound $\min_{x:~{}|x|=1}(|p(x)|/\sum_{i=0}^{d}|p_{i}|)\geq(\frac{\theta-1}{2\theta})^{d}$ of [134, Eqn. (9.4)], we would have only obtained a bound of order $m^{2}d$ on the computational precision of our root-finders.
@@ 17 Approximation of matrix eigenvalues
Given a $d\times d$ matrix $M$ that has characteristic polynomial $p(x)=\det(xI-M)$, our root-finders can approximate all $m$ its zeros (that is, the eigenvalues of $M$) lying in a fixed disc $D$ on the complex plane isolated from external eigenvalues.
Recall Gershgorin’s bound for any eigenvalue $\lambda$ of a matrix $M=(a_{i,j})_{i,j=1}^{d}$ [138, Thm. 1.3.2]: $|\lambda-a_{i,i}|\leq\sum_{j\neq i}|a_{i,j}|$ for some $i$, $1\leq i\leq d$.
@@
Combine this bound with the precision estimates of the previous section for our root-finders [^27] For a posteriori estimates, one can combine our root-finders with the recipe of doubling the precision of computing of (see Sec. 16.1). and with the following result of Storjohann [137] (cf. Remark 17.1):
@@
This implies a record bit operation complexity bound $\tilde{O}(m^{2}\tilde{b}^{2}d^{\omega})$ for any matrix $M$ which decreases whp by a factor of $m$ for matrices $M$ whose eigenvalues are iid random variables sampled under the uniform probability distribution from a large disc on the complex plane (cf. (1.6)) for approximation within $\epsilon$ of all $m$ eigenvalues of $M$ lying in the disc $D$ provided that $\tilde{b}=\log_{2}(||M||_{F}/\epsilon)$. This supports a bound exceeding (1.6) by a factor of $b$, to be removed with acceleration of our root-finding in Part III.
Remark 17.1.
Our root-finders are essentially reduced to the evaluation of NIR$(x)=\frac{p^{\prime}(x)}{p(x)}$ at sufficiently many points $x$, and we can rely on the following expressions for approximation of the eigenvalues of a matrix or a matrix polynomial:
$\frac{p(x)}{p^{\prime}(x)}=\frac{1}{{\rm trace}((xI-M)^{-1})}$
where $M$ is a matrix and $p(x)=\det(xI-M)$ (cf. (3.1)) and
$\frac{p(x)}{p^{\prime}(x)}=\frac{1}{{\rm trace}(M^{-1}(x)M^{\prime}(x))}$
where $M(x)$ is a matrix polynomial and $p(x)=\det(M(x))$ [20, Eqn. (5)]. [^28] These expressions have been obtained based on Jacobi’s formula ${\rm d}(\ln(\det(A)))={\rm trace}(A^{-1}{\rm d}A),$ which links differentials of a nonsingular matrix $A$ and of its determinant. Based on the above expressions one can reduce approximation of eigenvalues of a matrix or a matrix polynomial to its inversion. This computation tends to be more numerically stable than computing determinants and is significantly accelerated the highly important case where an input matrix or matrix polynomial can be inverted fast, e.g., is data sparse.
18 Bit operation complexity of general polynomial root-finding
Next we estimate the bit operation complexity of our root-finders, which accelerate classical subdivision root-finding by incorporating our e/i tests and are applied to a general polynomial $p$ of (1.1) whose coefficients are given with a precision sufficient to support our root-finding.
We could simply combine the precision bound $O(b\log(d))$ of Sec. 16 for our e/i tests with the estimate $O(bm)$ of Cor. 14.2 for the number of evaluation points for our root-finders, the bound of $2d$ ops for the evaluation of $p(x)$, and (12.5) for bit-operation cost of an op and then would arrive at the bit-operation complexity bound $\tilde{O}(b^{2}dm)$ for our root-finders. [^29] Recall that under the Random Root Model of Sec. 1.5 we can whp decrease this bound by a factor of $m$.
Next, however, we decrease this bound roughly by a factor of $d$ by using benefits of fast multipoint evaluation of a polynomial by means of the algorithm of Fiduccia [42] and Moenck and Borodin [84] (see [19, Sec. 4.5] for historical account), non-trivially extended in [81] to support the following theorem:

### v37

@@ 7.5 Counting the evaluation queries of our root-finders
Combine Cors. 2.2 and 7.1 and Prop. 7.4 and obtain
@@
We can reduce root-finding to deterministic evaluation of $p(x)$ at $O(m^{2}b\log(d))$ points as well as to the evaluation at $O(mb\log^{2}(d))$ points under the Random Root Model.
@@ 8.2 A priori estimates for computational precision
In the rest of this section we deduce a priori estimate of $O(b\log(d))$ bits for computational precision. Namely, we prove that this precision supports correctness of our subdivision root-finders; our analysis includes some advanced techniques of [59].
Estimating computational precision of our soft e/i tests we assume that they are applied to the unit disc $D(0,1)$ because we can extend the results to any disc $D(c,\rho)$ by applying map (1.4) and increasing the upper bounds on the precision by a factor of $\log_{2}(1/\rho)$. This factor is at most $b$ because we only handle discs of radii $\rho>\epsilon=1/2^{b}$ until we reach $\epsilon$-discs and stop. Therefore, our bound $O(\log(d))$ on the computational precision of our e/i tests for the disc $D(0,1)$ will imply the bound $O(b\log(d))$ for any disc.
We also make Assumption 0 that we can request and obtain from evaluation oracle (black box subroutine) the values of $p$ with a relative error bound within $1/d^{\nu}$ for any fixed constant $\nu>0$ of our choice; we can represent these values with the bit-precision $\log_{2}(d)+\log_{2}(\nu)$.
In our e/i tests we only need to compare the values $|{\rm NIR}_{t_{h}}(x)|$ at some complex points $x=v_{d}\geq 1/d^{\beta}$ for a constant $\beta>0$, and next we estimate that under Assumption 0 we can indeed do this by performing our computations with a precision of $O(\log(d))$ bits.
Multiplication with a precision $s$ contributes a relative error at most $2^{s}$ (cf. [14, Eqn. A.3 of Ch. 3]) and otherwise just sums the relative error bounds of the factors for we ignore the dominated impact of higher order errors. Hence we deduce that $h=O(\log(d))$ multiplications for root-lifting to their $h$th powers increase these bounds by at most a factor of $h$, which is immaterial since we allow relative errors of order $1/d^{\nu}$ for a constant $\nu>0$ of our choice.
We apply this analysis to approximation of NIR$(x)$ for $p_{h}(x)$ based on (3.2) and under the following
@@
This assumption holds whp because we apply our e/i tests under random rotation $x\mapsto vx$ for random variable $v$ specified in our next proposition.
@@
Let $\psi_{j}$ denote the length of the arc $C(0,1)\cap D(z_{j},\frac{2\pi}{mqw})$, $j=1,2,\dots,d$. Then $\psi_{j}=0$ for at least $d-m$ integers $j$, and $\max_{j=1}^{d}\psi_{j}\leq\frac{\pi^{2}}{mqw}$ for sure. Hence $\sum_{j=1}^{d}\psi_{j}<\frac{\pi^{2}}{qw}$, and so $\phi_{g}\leq\frac{2\pi}{mqw}$ for a fixed $g$ with a probability at most $\frac{\pi}{qw}$, and the lemma follows. ∎
Next, by applying Prop. 8.1 for a sufficiently large $w=d^{O(1)}$, we prove that a precision of $O(\log(d))$ bits supports our e/i tests under Assumptions 0 and 1. We will (i) prove that $|{\rm NIR}(z)|=d^{O(1)}$, (ii) will approximate this value within $d^{\nu}\delta$ for any fixed constant $\nu$ and for $\delta$ of (3.2) chosen such that $\log(\frac{1}{\delta})=O(\log(d))$, and (iii) will estimate that such a bound withstands the impact of rounding errors of computations for (3.2).
@@
In the proof of the next theorem we use the following lemma, which is [59, Fact 3.5].
@@
Under the assumptions of Prop. 8.2 it holds that
$|t(x)\alpha|=\Big{|}\frac{t(x)-t(x-\delta)}{\delta}-t^{\prime}(x)\Big{|}\leq\Big{|}t(x)\frac{(d\xi^{\prime})^{2}\delta}{1-d\xi^{\prime}\delta}\Big{|}~{}{\rm for}~{}\xi^{\prime}=\frac{1}{\rho}.$
@@
The claimed bound on $|\xi^{\prime}|$ follows from (3.1). It remains to apply the first five lines of the proof of [59, Lemma 3.6] with $f$, $\alpha$, $n$, and $\xi$ replaced by $t$, $-\delta$, $d$, and $w$, respectively. Namely, first obtain from Taylor’s expansion that $t(x)-t(x-\delta)=\sum_{j=0}^{\infty}\frac{\delta^{j}}{j!}t^{(j)}(x)-t(x)$.
@@
Suppose that the values $t(y)$ have been computed by an oracle within a relative error bounds $\nabla(y)$ for $y=x$, $y=x-\delta$, and $\delta$ of Cor. 8.2. Then one can ensure that $\beta<\frac{1}{d^{\bar{\eta}}}$ for any fixed constant $\bar{\eta}$ by choosing a proper $\nabla(y)$ of order $1/d^{O(1)}$, represented with a precision of $O(\log(d))$ bits, which we can do under Assumption 0.
@@
$\frac{t(x-\delta)}{t(x)}=1+\frac{t(x-\delta)-t(x)}{t(x)}=1+\delta\frac{t^{\prime}(u)}{t(x)},~{}{\rm for}~{}u\in[x-\delta,x].$
@@
By extending Cor. 8.2 obtain $|\alpha^{\prime}|\leq\frac{1}{d^{O(1)}}$ and deduce from (3.1) and Assumptions 0 and 1 that $|\frac{t^{\prime}(x)}{t(x)}|\leq\frac{1}{d^{O(1)}}$ (cf. the proof of Prop. 14.1). Hence
$w:=\Big{|}\frac{t(x-\delta)}{t(x)}\Big{|}\leq d^{O(1)}~{}{\rm and}~{}|\beta\delta|\leq|w|\frac{|\nabla(x)|+|\nabla(x-\delta)|}{1-|\nabla(x)|}.$
Now choose $\nabla(y)$ such that $|\nabla(y)|\leq\frac{|\delta|}{20|w|}$ for $y=x$ and $y=x-\delta$. Then verify that
$\log\Big{(}\frac{1}{|\nabla(y)|}\Big{)}=O(\log(d)),~{}1-|\nabla(x)|\geq 1-\frac{|\delta|}{20|w|}>\frac{19}{20},~{}{\rm and}~{}|\delta\beta|<\frac{|\delta|}{8}.$
Combine Cors. 8.1 and 8.2 and Prop. 8.4 to obtain
@@
Given an oracle for the evaluation of a polynomial $p(x)$ of a degree $d$ with a relative error of any order in $1/d^{O(1)}$ (this holds under Assumption 0), we can perform our black box subdivision root-finders with a precision of $O(\log(d))$ bits.
@@
We derived our precision bound $O(\log(d))$ (in the case of root-finding in the unit disc $D(0,1)$) based on operating entirely with NIR$(x)$. For comparison, if we relied on computing $p(x)$ and on the bound $\min_{x:~{}|x|=1}(|p(x)|/\sum_{i=0}^{d}|p_{i}|)\geq(\frac{\theta-1}{2\theta})^{d}$ of [106, Eqn. (9.4)], we would have only needed a bound of order $m^{2}d$ on the computational precision of our root-finders.
@@ 9 Approximation of matrix eigenvalues
Given a $d\times d$ matrix $M$ that has characteristic polynomial $p(x)=\det(xI-M)$, our root-finders can approximate all its $m$ zeros (that is, the eigenvalues of $M$) lying in a fixed disc $D$ on the complex plane isolated from its external zeros (or eigenvalues of $M$).
Recall Gershgorin’s bound for any eigenvalue $\lambda$ of a matrix $M=(a_{i,j})_{i,j=1}^{d}$ [111, Thm. 1.3.2]:
$|\lambda-a_{i,i}|\leq\sum_{j\neq i}|a_{i,j}|~{}{\rm for~{}some}~{}i,~{}1\leq i\leq d.$
@@
Combine this bound with the a priori precision estimates of the previous section for our root-finders [^16] For a posteriori estimates for the bit operation complexity of Problems 0, $0^{*}$, 2, and 2, one can combine our root-finders with the recipe of doubling the precision of computing of Sec. 8.1. and with the following result of Storjohann [109]:
@@
For any $d\times d$ matrix $M$ we arrive at a record bit operation complexity bound $\tilde{O}(m^{2}\tilde{b}bd^{\omega})$ for approximation within $\epsilon$ of all $m$ eigenvalues of any matrix $M$ lying in the disc $D$ provided that $\tilde{b}=\log_{2}(||M||_{F}/\epsilon)$. This bound exceeds (1.6) by a factor of $b$ and can be decreased by a factor of $m$ under the Random Root Model for the roots of the characteristic polynomial of $M$.
10 Bit operation complexity of polynomial root-finding
10.1 Straightforward extension of the precision bound and a direction to saving evaluation queries
Recall the bound
$\mu(s):=O(s\log(s)=\tilde{O}(s)$ (10.1)
on the number of bit operations for integer multiplication modulo $2^{s}$ [40] (cf. [114, 53, 106, 107]), combine it with the estimate $O(bm)$ of Cor. 7.2 for the number of evaluation queries in our root-finders, the upper bound $2d$ ops for the evaluation of a polynomial $p(x)$ given with its coefficients, and the precision bound $O(b\log(d))$ of Sec. 8, and obtain the bit-operation complexity bound $\tilde{O}(b^{2}dm)$ for our root-finders. [^17] By extending precision bound of [59] we greatly simplified derivation of our bit-complexity estimates versus the papers [106, 52], each of 46 pages, which bypass precision bound and had to use long tedious step-by-step analysis of their much involved root-finders.
Since $p(x)$ is given with its coefficients, we apply fast multipoint evaluation of a polynomial in [32, 67] and use fewer ops, roughly by a factor of $d/\log^{2}(d)$. Peter Kirrinnis in [52] and Moroz in [64] have non-trivially extended this progress to saving bit-operations: